# Celestijnenlaan 200D, B-3001 Leuven, Belgium

### v1

@@ 3.3 Euclidean on-shell actions
We would like to highlight that, to the best of our knowledge, this represents the first rigorous calculation of the Euclidean on-shell action for electrically charged, rotating AdS7 black holes in 7d maximal gauged supergravity, achieved through proper holographic renormalization. In previous studies, even for special cases with a reduced parameter space – such as identical rotations Chong:2004dy , identical charges Chow:2007ts , and single rotation Wu:2011gp – the regularized on-shell action was computed indirectly, relying on the quantum statistical relation between the conserved charges of the black hole background and the corresponding chemical potentials, see also Cassani:2019mms . Hence, our calculation places a holographic comparison between the regularized on-shell action of the Euclidean BDHM background and the dual $S^{1}\times S^{5}$ superconformal index (SCI) of the 6d $\mathcal{N}=(2,0)$ theory on a more robust footing, without resorting to black hole thermodynamics. In what follows, we discuss the regularized on-shell action of the BDHM background (48) in this context of holography, with particular emphasis on the choice of finite counterterms.
@@ 3.4 Holography
The AdS/CFT correspondence relates the Euclidean M-theory path integral around the 11d supergravity background obtained by uplifting the supersymmetric limit of the BDHM solution via the consistent truncation of Nastase:1999cb ; Nastase:1999kf ; Liu:1999ai to the dual SCI defined as supersymmetric path integral of the $\mathcal{N}=(2,0)$ $A_{N}$ SCFT on $S^{1}\times S^{5}$ Choi:2018hmj ; Hosseini:2018dob ; Cassani:2019mms ; Nahmgoong:2019hko ; Bobev:2015kza ; Ohmori:2020wpk ; Bobev:2023bxl . In the limit where the AdS radius is much larger than the 11d Planck length, the M-theory path integral is approximated by the regularized on-shell action of the supersymmetric Euclidean BDHM background in 7d gauged supergravity. According to the AdS/CFT correspondence, this regularized on-shell action should agree with the large $N$ limit of the dual SCI. This relation was explicitly verified in Bobev:2023bxl provided that
@@
Another important point is that applying our choice of finite counterterms to the global EAdS7 background, i.e. empty AdS7, yields a vanishing regularized on-shell action. Therefore, it does not resolve the mismatch issue in the AdS7/CFT6 correspondence regarding the $\frac{81}{80}$ factor of discrepancy between the regularized on-shell action of global EAdS7 and the dual field theory quantity encoded in the $S^{5}$ free energy of 5d SYM, as discussed in Kallen:2012zn ; Minahan:2013jwa ; Bobev:2019bvq . The standard local counterterms introduced in subsection 3.1 are insufficient to resolve this mismatch, and one may instead consider giving up gauge invariance of the counterterms as suggested in Bobev:2019bvq . Thus, despite the progress we make here, a unified understanding of holographic renormalization, including finite counterterms, for all asymptotically locally EAdS7 backgrounds remains a subject for future work.
@@ 4.3.2 Supersymmetry
Following the supersymmetry analysis of Cvetic:2005zi , as applied in Chow:2007ts ; Cassani:2019mms ; Bobev:2023bxl , we restrict our attention to the supersymmetric limit preserving two real supercharges, imposed by the following constraint on the black hole parameters:
@@
The supersymmetric regularized Euclidean on-shell action (80) is expected to capture, in the semi-classical limit, the supersymmetric path integral of the dual 6d $\mathcal{N}=(2,0)$ theory over $S^{1}\times L^{p,q,r}$ on the ‘2nd sheet’. This result generalizes prior holographic comparisons for the 6d SCFT on $S^{1}\times S^{5}$ Hosseini:2018dob ; Choi:2018hmj ; Cassani:2019mms ; Bobev:2023bxl . We now proceed to elaborate on this holographic interpretation.
@@ 5 AdS5 black holes with generalized conformal boundaries
We focus on the 5d gauged supergravity coupled to two vector multiplets, referred to as the gauged STU model. In this theory, the most general black hole solution with two rotation parameters and three electric charge parameters was constructed in Wu:2011gq . For the purpose of generalization here, we focus on a special case where the rotation parameters are identical, which was first found in Cvetic:2004ny and further studied in Cvetic:2005zi ; Cassani:2019mms . Following the conventions of Cvetic:2005zi ; Cassani:2019mms , the solution reads
@@ Acknowledgments
We are grateful to Pieter Bomans, Davide Cassani, Friðrik Freyr Gautason, Dario Martelli, and Rishi Mouland for useful discussions. NB is supported in part by FWO projects G003523N, G094523N, and G0E2723N, as well as by the Odysseus grant G0F9516N from the FWO. JH is supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) with grant number RS-2024-00449284, the Sogang University Research Grant of 202410008.01, the Basic Science Research Program of the National Research Foundation of Korea (NRF) funded by the Ministry of Education through the Center for Quantum Spacetime (CQUeST) with grant number RS-2020-NR049598. MD is supported by the Postdoctoral Fellowship of the Research Foundation- Flanders with the grant 1235324N. VD is partially supported by a grant Trapezio (2023) of the Fondazione Compagnia di San Paolo and would like to thank the ITF at KU Leuven for hospitality during various stages of this project.

### v2

@@ 3.3 Euclidean on-shell actions
We would like to highlight that, to the best of our knowledge, this represents the first rigorous calculation of the Euclidean on-shell action for electrically charged, rotating AdS7 black holes in 7d maximal gauged supergravity, achieved through proper holographic renormalization. In previous studies, even for special cases with a reduced parameter space – such as identical rotations Chong:2004dy, identical charges Chow:2007ts, and single rotation Wu:2011gp – the regularized on-shell action was computed indirectly, relying on the quantum statistical relation between the conserved charges of the black hole background and the corresponding chemical potentials, see also Cassani:2019mms; Kantor:2019lfo. Hence, our calculation places a holographic comparison between the regularized on-shell action of the Euclidean BDHM background and the dual $S^{1}\times S^{5}$ superconformal index (SCI) of the 6d $\mathcal{N}=(2,0)$ theory on a more robust footing, without resorting to black hole thermodynamics. In what follows, we discuss the regularized on-shell action of the BDHM background (48) in this context of holography, with particular emphasis on the choice of finite counterterms.
@@ 3.4 Holography
The AdS/CFT correspondence relates the Euclidean M-theory path integral around the 11d supergravity background obtained by uplifting the supersymmetric limit of the BDHM solution via the consistent truncation of Nastase:1999cb; Nastase:1999kf; Liu:1999ai to the dual SCI defined as supersymmetric path integral of the $\mathcal{N}=(2,0)$ $A_{N}$ SCFT on $S^{1}\times S^{5}$ Choi:2018hmj; Hosseini:2018dob; Cassani:2019mms; Kantor:2019lfo; Nahmgoong:2019hko; Bobev:2015kza; Ohmori:2020wpk; Bobev:2023bxl. In the limit where the AdS radius is much larger than the 11d Planck length, the M-theory path integral is approximated by the regularized on-shell action of the supersymmetric Euclidean BDHM background in 7d gauged supergravity. According to the AdS/CFT correspondence, this regularized on-shell action should agree with the large $N$ limit of the dual SCI. This relation was explicitly verified in Bobev:2023bxl provided that
@@
Another important point is that applying our choice of finite counterterms to the global EAdS7 background, i.e. empty AdS7, yields a vanishing regularized on-shell action. Therefore, it does not resolve the mismatch issue in the AdS7/CFT6 correspondence regarding the $\frac{81}{80}$ factor of discrepancy between the regularized on-shell action of global EAdS7 and the dual field theory quantity, which is encoded in the $S^{5}$ free energy of 5d SYM through the perspective originally proposed in Douglas:2010iu; Lambert:2010iw, as discussed in Kallen:2012zn; Minahan:2013jwa; Bobev:2019bvq. The standard local counterterms introduced in subsection 3.1 are insufficient to resolve this mismatch, and one may instead consider giving up gauge invariance of the counterterms as suggested in Bobev:2019bvq. Thus, despite the progress we make here, a unified understanding of holographic renormalization, including finite counterterms, for all asymptotically locally EAdS7 backgrounds remains a subject for future work.
@@ 4.3.2 Supersymmetry
Following the supersymmetry analysis of Cvetic:2005zi, as applied in Chow:2007ts; Cassani:2019mms; Kantor:2019lfo; Bobev:2023bxl, we restrict our attention to the supersymmetric limit preserving two real supercharges, imposed by the following constraint on the black hole parameters:
@@
The supersymmetric regularized Euclidean on-shell action (80) is expected to capture, in the semi-classical limit, the supersymmetric path integral of the dual 6d $\mathcal{N}=(2,0)$ theory over $S^{1}\times L^{p,q,r}$ on the ‘2nd sheet’. This result generalizes prior holographic comparisons for the 6d SCFT on $S^{1}\times S^{5}$ Hosseini:2018dob; Choi:2018hmj; Cassani:2019mms; Kantor:2019lfo; Bobev:2023bxl. We now proceed to elaborate on this holographic interpretation.
@@ 5 AdS5 black holes with generalized conformal boundaries
We focus on the 5d gauged supergravity coupled to two vector multiplets, referred to as the gauged STU model. In this theory, the most general black hole solution with two rotation parameters and three electric charge parameters was constructed in Wu:2011gq. For the purpose of generalization here, we focus on a special case where the rotation parameters are identical, which was first found in Cvetic:2004ny and further studied in Cvetic:2005zi; Cassani:2019mms; Kantor:2019lfo. Following the conventions of Cvetic:2005zi; Cassani:2019mms, the solution reads
@@ Acknowledgments
We are grateful to Pieter Bomans, Davide Cassani, Friðrik Freyr Gautason, Robie Hennigar, Dario Martelli, and Rishi Mouland for useful discussions. NB is supported in part by FWO projects G003523N, G094523N, and G0E2723N, as well as by the Odysseus grant G0F9516N from the FWO. JH is supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) with grant number RS-2024-00449284, the Sogang University Research Grant of 202410008.01, the Basic Science Research Program of the National Research Foundation of Korea (NRF) funded by the Ministry of Education through the Center for Quantum Spacetime (CQUeST) with grant number RS-2020-NR049598. MD is supported by the Postdoctoral Fellowship of the Research Foundation- Flanders with the grant 1235324N. VD is partially supported by a grant Trapezio (2023) of the Fondazione Compagnia di San Paolo and would like to thank the ITF at KU Leuven for hospitality during various stages of this project.